Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 601     Entry time: 02 Apr 2025, 10:31
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 02.04.2025 
General Announcement:
- A. Galler reminds instruments that enquiring and requests on beam parameters shall come prior the respective RC meeting. Otherwise it will be difficult to accommodate last minute 
changes.

CTRL:
- Performed update of dataloggers yesterday.

DA:
- All fine with DA.

EEE:
- Nothing to report.

VAC:
- Issue with broken power supply in SA3 on Monday 2PM. Downtime of 2 hs and it was fixed by exchanging it in a ZZ.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Not present.

LAS:
- Laser system working;
- Last Thursday evening decisions were done that incurred effects on PP-laser and access. An e-mail to XO will be sent with details.

UNSYS:
- Everything is fine.

SPB:
- Nothing to report.

FXE:
- Reported motor issues are fixed.

MID:
- Nothing to report.

HED:
- Not present.

SCS:
- Nothing to report; busy preparing DSSC2 setup. 

SQS:
- Nothing to report; ready for beam.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00